# The Need for a Socially-Grounded Persona Framework for User Simulation

### v1

@@ Abstract
Synthetic personas are widely used to condition large language models (LLMs) for social simulation, yet most personas are still constructed from coarse sociodemographic attributes or summaries. We revisit persona creation by introducing SCOPE, a socially grounded framework for persona construction and evaluation, built from a 141-item, two-hour sociopsychological protocol collected from 124 U.S.-based participants. Across seven models, we find that demographic-only personas are a structural bottleneck: demographics explain only $\sim$1.5% of variance in human response similarity. Adding sociopsychological facets improves behavioral prediction and reduces over-accentuation, and non-demographic personas based on values and identity achieve strong alignment with substantially lower bias. These trends generalize to SimBench (441 aligned questions), where SCOPE personas outperform default prompting and NVIDIA Nemotron personas, and SCOPE augmentation improves Nemotron-based personas. Our results indicate that persona quality depends on sociopsychological structure rather than demographic templates or summaries.
@@ 1 Introduction
Figure 1: End-to-end pipeline of our persona construction and evaluation framework. The process consists of four major stages: (1) Rich Persona Collection: a two-hour, 141-item sociopsychological survey capturing eight behavioral facets across 124 participants; (2) Persona Construction Cases: structured persona representations generated by selectively combining demographic, behavioral, psychological, and narrative facets, as well as AI-generated summaries; (3) Synthetic Persona Evaluation: measuring accuracy, similarity to human responses, and demographic accentuation; and
@@ 2.2 Persona Construction in NLP
Identity Narratives $\bullet$ $\times$ $\bullet$ $\bullet$ $\bullet$
Professional Identity $\times$ $\times$ $\times$ $\circ$ $\bullet$
@@ 3 Socially Grounded Persona Framework
Synthetic personas are widely used in NLP and AI, yet their construction remains under-theorized and weakly grounded in empirical social science. Prior work in computational personality and behavior modeling shows that psychological traits and values are substantially more predictive of behavior than sociodemographic variables alone, a distinction largely absent from current persona construction practices Ketipov et al. (2023); Joshi et al. (2025). To address this gap, we introduce SCOPE: Sociopsychological Construct of Persona Evaluation, a sociopsychologically informed framework for constructing rich, multidimensional persona representations.
@@ 3.2 Data Collection Design
Each facet was informed by established social-science frameworks. Personality items followed the BFI-2-S, value items drew on Schwartz’s value theory, and behavioral items were aligned with large-scale surveys such as Pew Internet & Technology, the General Social Survey (GSS), and the World Values Survey (WVS) Smith et al. (2023); Pew Research Center (2024); Inglehart et al. (2014). All items were written as an original composite survey rather than directly reproducing any single instrument. We recruited a U.S.-based sample guided by Census-like proportions, with explicit emphasis on racial diversity. Recruitment occurred in two stages: (1) outreach via Twitter/X, LinkedIn, and professional Slack groups; and (2) stratified sampling through a commercial research platform (User Interviews). The study protocol was approved by the institution’s Ethics Office.
@@ 3.3 Synthetic Persona Construction
Case 7: Non-Demographic Personas. Personas that omit demographics entirely (e.g., identity-only or values+identity personas). These variants evaluate whether non-demographic cues alone can support behavioral simulation.
@@ 4.1 Baseline Persona Comparisons
II. Adding sociopsychological facets improves alignment. Sequentially enriching personas with identity narratives (Case 2), values (Case 3), and full SCOPE conditioning (Case 4) steadily increases correlation and accuracy, with Case 4 reaching $\bar{r}=0.667$ and 39.7% accuracy.
III. Human-grounded facets outperform AI-compressed personas. AI-generated narrative summaries (Case 5c) slightly improve over demographics but never surpass fully grounded representations, even when summaries are long. Non-demographic SCOPE personas built only from values and identity (Case 7c) match or exceed the best full-information case in correlation while avoiding direct use of demographic attributes.
IV. Personas grounded in non-sociodemographic information perform better. The greater finding show that to understand social behaviours, sociodemographic information is not the most important element of a persona. We see that for Case 6 iterations, where personas are created using either or both values or identity information tend to perform better, in regards to accuracy, than the ones generated by sociodemographic information.
@@
2: + Identity narratives 0.642 37.74 -8.30
3: + Values & motivations 0.656 38.60 75.50
@@
6f: AI identity+values 0.608 43.42 -40.40
7c: Values + identity only 0.658 39.68 -56.35
@@ 4.3 Bias and Behavioral Disparities
Second, adding sociopsychological grounding substantially reduces demographic accentuation, though the magnitude varies by facet. For GPT-4o, Case 2 (demographics + identity) reduces bias to near-baseline ($\mathrm{Bias\%}=8.30$), whereas Case 3 (demographics + values) still shows non-trivial over-accentuation ($\mathrm{Bias\%}=75.50$), suggesting that values alone do not necessarily prevent demographic clustering when demographics remain explicit inputs. In contrast, Case 4 (full conditioning) slightly under-shoots the human baseline for GPT-4o ($\mathrm{Bias\%}=-6.40$), indicating that SCOPE can preserve behavioral alignment.
Third, non-demographic personas are the most conservative with respect to demographic bias. For GPT-4o, Case 7c (values + identity, no demographics) yields $\mathrm{Bias\%}=-56.35$ while maintaining high correlation with human response structure ($\bar{r}=0.658$; Table 3). A similar pattern appears across other models in Appendix Tables 1617, where identity- and values-driven variants tend to be negative (or substantially reduced) in Bias% relative to demographic-only variants.
@@ 4.4 Group-based SCOPE Persona Analysis
Incorporating sociopsychological information mitigates this effect. Case 2 (Demographics + Identity) and Case 3 (Demographics + Values) reduce bias across most groups. In contrast, Case 4 (Full Conditioning) brings AI behavior close to the human baseline for several races and yields near-zero or negative bias for African American or Black and Hispanic or Latino participants. Ideally we see that Demographics + Values + Augmented show the most ideal required alignment without stereotyping or misrepresentation. Negative bias can also be a problem where the persona created is too dissimilar to the group in question. Therefore, a balance is needed for the right persona behaviour without stereotyping or misrepresenting a group.
@@
Case_2: Demographics + Identity +0.070 0.113
Case_3: Demographics + Values +0.165 0.028
@@
Case_7a: Identity Only -0.049 0.145
Case_7b: Values Only -0.068 0.102
Case_7c: Identity + Values -0.053 0.088
@@
Table 8 summarizes bias accentuation averaged across racial groups. Demographics-only personas (Case 1 and Case 1S) show the highest average bias, confirming that reliance on demographic attributes alone encourages stereotyping. Hybrid personas that combine demographics with identity or values reduce bias but remain positive on average. In contrast, personas that remove explicit demographic information, such as identity-only, values-only, or augmented non-demographic variants, consistently yield negative bias, indicating reduced demographic dependence relative to humans but also possible misrepresentation of a group. We also see in cases without demographic information, there is complete loss in patterns of behacioural ques leading to negative bias as well. These results show that demographic based group bias in persona-based simulations is neither uniform nor inevitable. Bias emerges most strongly for specific racial groups under specific persona construction strategies, and can be systematically reduced through sociopsychological grounding. We analyze just race in this subsection but each sociodemographic attribute provides new insights on the dependency of the model generating the persona.
@@ 5 Discussion and Conclusion
Sociopsychological grounding improves alignment. Adding values, traits, identity narratives, and behavioral signals steadily improves alignment across models, with full conditioning (Case 4) achieving the highest correlation while reducing demographic bias. These gains are not driven by verbosity: AI-generated summaries (Case 5) are often longer yet underperform structured, human-grounded representations, indicating that persona structure matters as much as information quantity. Consistently, non-demographic personas based on values and identity (Case 7c) match or exceed full conditioning in correlation while substantially reducing bias, showing that demographics are neither necessary nor sufficient.
Bias emerges from persona design choices. Demographic-only personas systematically over-accentuate demographic clustering, effectively encoding stereotypes into simulated behavior. In contrast, SCOPE reduces bias without sacrificing realism: full conditioning closely matches the human baseline, while values+identity personas are the most conservative with respect to demographic amplification while maintaining strong alignment. This supports a practical principle: bias reduction is best achieved by replacing coarse demographic proxies with psychologically meaningful structure, rather than removing structure altogether.
@@
We collected a two-hour sociopsychological questionnaire from U.S.-based adult participants under review by the institution’s Ethics Office. Participants provided informed consent and were compensated \$50 for completion. To reduce privacy risk, we avoided collecting and removed direct identifiers (e.g., legal names, email addresses, phone numbers, street addresses). Free-text responses were stored under anonymized participant IDs and screened to remove accidental disclosure of sensitive personal information. Because persona conditioning can amplify stereotypes, we explicitly measure demographic accentuation (Section 4.3) and report persona variants that achieve strong behavioral alignment while reducing demographic clustering. We plan to release only the Nemotron augmented data, the pipeline code to create augmented SCOPE personas, AI profile materials, and report aggregate statistics to omit any re-identification risk. We will release the GitHub repository soon following internal review process and cleanup. In the interim, interested researchers are welcome to contact the authors; data access can be considered on a case-by-case basis subject to appropriate review.
@@
No sensitive protected health information or financial identifiers were collected. The anonymized dataset used for modeling contains only aggregated demographic categories and de-identified narrative text. Access to the full dataset is restricted to the research team, and all analyses reported in this paper were conducted on the de-identified corpus. These procedures were designed to provide strong privacy protection while enabling rigorous sociotechnical research on persona-grounded AI systems. In regards to AI usage in the paper, we used AI-assisted tools solely for grammatical and stylistic corrections of our writing. All scientific content, analysis, and conclusions are the authors’ own.
@@ Appendix A Appendix
Refs: Park et al. (2023), Pew Research Center (2024), Smith et al. (2023), Inglehart et al. (2014)
@@ A.7 External Data Formats: Nemotron Personas and SimBench Instances
Case 3: Demographics + Values & Goals Demographics plus structured value items and goal statements (Likert-style), emphasizing motivational priorities.
@@
Values/Goals (snippet): “Helping people around me”=5; “Being rich”=2; “Equal opportunities”=5; “Taking risks/adventures”=3; “Secure surroundings”=4; “Thinking up new ideas/being creative”=4.
@@
Demographics + Values/Goals
Case 4: Full Conditioning Combined structured persona: demographics + values/goals + professional identity + personal identity narratives.
@@
Values/Goals (snippet): “Environmental protection”=5; “Tradition”=2; “Achievement/recognition”=3; “Independence”=5.
@@
Demographics + Values/Goals + Professional Identity + Personal Identity
@@
Case 7: Non-Demographic Persona Persona excludes demographics entirely; conditioning uses non-demographic signals (e.g., identity-only, values-only, or values+identity).
Example (values+identity only):
@@
Values/Goals (snippet): “Equality”=5; “Loyalty to friends”=5; “Tradition”=2; “Independence”=4; “Creativity”=3.
@@
Values and/or Identity (no demographics)
@@
Case_2 Demographics + Identity 0.642 37.74 0.011 8.298 0.656 38.56 0.012 8.791 0.609 33.58 -0.101 -76.79 0.616 27.58 -0.016 -12.245
Case_3 Demographics + Values 0.656 38.60 0.100 75.499 0.650 39.60 0.055 41.287 0.600 36.71 0.012 9.087 0.614 22.56 -0.042 -31.64
@@
Case_6b AI (Given Demography + Identity) 0.619 25.55 -0.034 -25.81 0.435 30.18 0.029 21.698 0.649 20.01 -0.041 -30.71 0.517 25.43 -0.065 -49.244
Case_6c AI (Given Demography + Values) 0.606 27.86 0.016 12.249 0.462 30.15 0.009 6.797 0.681 21.32 -0.025 -19.097 0.496 27.30 -0.066 -50.171
@@
Case_6e AI (Given Values alone) 0.588 41.65 -0.080 -71.43 0.592 41.45 -0.071 -63.02 0.544 40.49 -0.101 -90.04 0.577 40.37 -0.099 -87.90
Case_6f AI (Given Identity + Values) 0.608 43.42 -0.046 -40.40 0.614 43.15 -0.038 -33.91 0.561 40.76 -0.079 -70.22 0.586 42.43 -0.082 -72.79
@@
Case_7b Values Only 0.624 37.95 -0.094 -71.092 0.639 38.92 -0.088 -66.353 0.595 36.28 -0.121 -91.436 0.595 34.27 -0.115 -87.329
Case_7c Values+Identity 0.658 39.68 -0.074 -56.354 0.662 40.29 -0.053 -40.485 0.612 37.86 -0.099 -74.563 0.636 33.89 -0.101 -76.371
@@
Case_2 Demographics + Identity 0.609 35.40 -0.022 -0.207 0.615 31.80 0.035 0.322 0.572 20.90 -0.020 -0.190
Case_3 Demographics + Values 0.602 35.10 -0.011 -0.107 0.591 29.10 0.050 0.469 0.553 19.10 -0.033 -0.304
@@
Case_6c AI (Given Demography + Values) 0.560 25.76 -0.018 -13.811 0.564 23.22 -0.015 -11.300 0.521 15.79 -0.017 -13.183
Case_6d AI (Given Identity alone) 0.528 36.73 -0.082 -72.564 0.532 33.10 -0.067 -59.371 0.492 22.51 -0.078 -69.266
Case_6e AI (Given Values alone) 0.543 38.52 -0.097 -85.907 0.547 34.71 -0.079 -70.288 0.506 23.61 -0.092 -82.002
Case_6f AI (Given Identity + Values) 0.561 40.15 -0.067 -59.763 0.566 36.19 -0.055 -48.897 0.523 24.61 -0.064 -57.046
Case_7a Identity Only 0.584 34.62 -0.105 -79.570 0.589 31.20 -0.086 -65.103 0.544 21.22 -0.100 -75.953
Case_7b Values Only 0.576 35.10 -0.115 -86.958 0.581 31.63 -0.094 -71.147 0.537 21.51 -0.110 -83.005
Case_7c Values+Identity 0.608 36.70 -0.090 -68.138 0.612 33.07 -0.074 -55.749 0.566 22.49 -0.086 -65.040

### v2

@@ Abstract
Synthetic personas are widely used to condition large language models (LLMs) for social simulation, yet most personas are still constructed from coarse sociodemographic attributes or summaries. We revisit persona creation by introducing SCOPE, a socially grounded framework for persona construction and evaluation, built from a 141-item, two-hour sociopsychological protocol collected from 124 U.S.-based participants. Across seven models, we find that demographic-only personas are a structural bottleneck: demographics explain only $\sim$1.5% of variance in human response similarity. Adding sociopsychological facets improves behavioral prediction and reduces over-accentuation, and non-demographic personas based on traits and identity achieve strong alignment with substantially lower bias. These trends generalize to SimBench (441 aligned questions), where SCOPE personas outperform default prompting and NVIDIA Nemotron personas, and SCOPE augmentation improves Nemotron-based personas. Our results indicate that persona quality depends on sociopsychological structure rather than demographic templates or summaries.
@@ 1 Introduction
Figure 1: End-to-end pipeline of our persona construction and evaluation framework. The three major stages are: (1) Rich Persona Collection: a two-hour, 141-item sociopsychological survey capturing eight facets across 124 participants; (2) Persona Construction Cases: structured persona representations generated by selectively combining demographic, behavioral, psychological, and narrative facets, as well as AI-generated summaries; (3) Synthetic Persona Evaluation: measuring accuracy, similarity to human responses, and group bias.
@@ 2.2 Persona Construction in NLP
Identity Narratives $\bullet$ $\circ$ $\bullet$ $\bullet$ $\bullet$
Professional Identity $\times$ $\bullet$ $\times$ $\circ$ $\bullet$
@@ 3 Socially Grounded Persona Framework
Synthetic personas are widely used in NLP and AI, yet their construction remains under-theorized and weakly grounded in empirical social science. Prior work in computational personality and behavior modeling shows that psychological traits and values are substantially more predictive of behavior than sociodemographic variables alone, a distinction largely absent from current persona construction practices Ketipov et al. (2023); Joshi et al. (2025). To address this gap, we introduce SCOPE: Sociopsychological Construct of Persona Evaluation, a sociopsychologically informed framework for constructing rich, multidimensional persona.
@@ 3.2 Data Collection Design
Each facet was informed by established social-science frameworks. Personality items followed the BFI-2-S, value items drew on Schwartz’s value theory, and behavioral items were aligned with large-scale surveys such as Pew Internet & Technology, the General Social Survey (GSS), and the World Values Survey (WVS) Davis and Smith (1999); Pew Research Center (2024); Inglehart et al. (2014). All items were written as an original composite survey rather than directly reproducing any single instrument. We recruited a U.S.-based sample guided by Census-like proportions, with explicit emphasis on racial diversity. Recruitment occurred in two stages: (1) outreach via Twitter/X, LinkedIn, and professional Slack groups; and (2) stratified sampling through a commercial research platform (User Interviews). The study protocol was approved by the institution’s Ethics Office.
@@ 3.3 Synthetic Persona Construction
Case 7: Non-Demographic Personas. Personas that omit demographics entirely (e.g., narrative-only or trait+narrative personas). These variants evaluate whether non-demographic cues alone can support behavioral simulation.
@@ 4.1 Baseline Persona Comparisons
II. Adding sociopsychological facets improves alignment. Sequentially enriching personas with identity narratives (Case 2), traits (Case 3), and full SCOPE conditioning (Case 4) steadily increases correlation and accuracy, with Case 4 reaching $\bar{r}=0.667$ and 39.7% accuracy.
III. Human-grounded facets outperform AI-compressed personas. AI-generated narrative summaries (Case 5c) slightly improve over demographics but never surpass fully grounded representations, even when summaries are long. Non-demographic SCOPE personas built only from traits and idenitity narratives (Case 7c) match or exceed the best full-information case in correlation while avoiding direct use of demographic attributes.
IV. Personas grounded in non-sociodemographic information perform better. The greater finding show that to understand social behaviours, sociodemographic information is not the most important element of a persona. We see that for Case 6 iterations, where personas are created using either or both traits or identity narrative information tend to perform better, in regards to accuracy, than the ones generated by sociodemographic information.
@@
2: + Narratives 0.642 37.74 -8.30
3: + Traits 0.656 38.60 75.50
@@
6f: AI narrative+traits 0.608 43.42 -40.40
7c: traits + narrative only 0.658 39.68 -56.35
@@ 4.3 Bias and Behavioral Disparities
Second, adding sociopsychological grounding substantially reduces demographic accentuation, though the magnitude varies by facet. For GPT-4o, Case 2 (demographics + identity) reduces bias to near-baseline ($\mathrm{Bias\%}=8.30$), whereas Case 3 (demographics + traits) still shows non-trivial over-accentuation ($\mathrm{Bias\%}=75.50$), suggesting that traits alone do not necessarily prevent demographic clustering when demographics remain explicit inputs. In contrast, Case 4 (full conditioning) slightly under-shoots the human baseline for GPT-4o ($\mathrm{Bias\%}=-6.40$), indicating that SCOPE can preserve behavioral alignment.
Third, non-demographic personas are the most conservative with respect to demographic bias. For GPT-4o, Case 7c (traits + identity, no demographics) yields $\mathrm{Bias\%}=-56.35$ while maintaining high correlation with human response structure ($\bar{r}=0.658$; Table 3). A similar pattern appears across other models in Appendix Tables 1617, where identity- and traits-driven variants tend to be negative (or substantially reduced) in Bias% relative to demographic-only variants.
@@ 4.4 Group-based SCOPE Persona Analysis
Incorporating sociopsychological information mitigates this effect. Case 2 (Demographics + Narrative) and Case 3 (Demographics + Traits) reduce bias across most groups. In contrast, Case 4 (Full Conditioning) brings AI behavior close to the human baseline for several races and yields near-zero or negative bias for African American or Black and Hispanic or Latino participants. Ideally we see that Demographics + Traits + Augmented show the most ideal required alignment without stereotyping or misrepresentation. Negative bias can also be a problem where the persona created is too dissimilar to the group in question. Therefore, a balance is needed for the right persona behaviour without stereotyping or misrepresenting a group.
@@
Case_2: Demographics + Narratives +0.070 0.113
Case_3: Demographics + Traits +0.165 0.028
@@
Case_7a: Narratives Only -0.049 0.145
Case_7b: Traits Only -0.068 0.102
Case_7c: Narratives + Traits -0.053 0.088
@@
Table 8 summarizes bias accentuation averaged across racial groups. Demographics-only personas (Case 1 and Case 1S) show the highest average bias, confirming that reliance on demographic attributes alone encourages stereotyping. Hybrid personas that combine demographics with narratives or traits reduce bias but remain positive on average. In contrast, personas that remove explicit demographic information, such as identity-only, traits-only, or augmented non-demographic variants, consistently yield negative bias, indicating reduced demographic dependence relative to humans but also possible misrepresentation of a group. We also see in cases without demographic information, there is complete loss in patterns of behacioural ques leading to negative bias as well. These results show that demographic based group bias in persona-based simulations is neither uniform nor inevitable. Bias emerges most strongly for specific racial groups under specific persona construction strategies, and can be systematically reduced through sociopsychological grounding. We analyze just race in this subsection but each sociodemographic attribute provides new insights on the dependency of the model generating the persona.
@@ 5 Discussion and Conclusion
Sociopsychological grounding improves alignment. Adding values, traits, identity narratives, and behavioral signals steadily improves alignment across models, with full conditioning (Case 4) achieving the highest correlation while reducing demographic bias. These gains are not driven by verbosity: AI-generated summaries (Case 5) are often longer yet underperform structured, human-grounded representations, indicating that persona structure matters as much as information quantity. Consistently, non-demographic personas based on traits and narrative identity (Case 7c) match or exceed full conditioning in correlation while substantially reducing bias, showing that demographics are neither necessary nor sufficient.
Bias emerges from persona design choices. Demographic-only personas systematically over-accentuate demographic clustering, effectively encoding stereotypes into simulated behavior. In contrast, SCOPE reduces bias without sacrificing realism: full conditioning closely matches the human baseline, while traits+narrative personas are the most conservative with respect to demographic amplification while maintaining strong alignment. This supports a practical principle: bias reduction is best achieved by replacing coarse demographic proxies with psychologically meaningful structure, rather than removing structure altogether.
@@
We release only the SCOPE personas, the pipeline code to create augmented SCOPE personas, AI profile materials, and report aggregate statistics to omit any re-identification risk. We release the evaluation code and the SCOPE persona framework in the following Github repository [^5] https://github.com/SalesforceAIResearch/SCOPE-Persona-Framework/ from which you can generate your own SCOPE augmented personas given a set of seed information. We also release our set of SCOPE persona collection in Hugging Face [^6] https://huggingface.co/datasets/Salesforce/SCOPE-Persona/, which consists of 1M synthetic personas that are curated to have additional faceted social and behavioural information. These personas have shown to outperform the existing persona frameworks, especially in behavioural and social tasks, and hence can be used in similar environment settings.
@@
We collected a two-hour sociopsychological questionnaire from U.S.-based adult participants under review by the institution’s Ethics Office. Participants provided informed consent and were compensated \$50 for completion. To reduce privacy risk, we avoided collecting and removed direct identifiers (e.g., legal names, email addresses, phone numbers, street addresses). Free-text responses were stored under anonymized participant IDs and screened to remove accidental disclosure of sensitive personal information. Because persona conditioning can amplify stereotypes, we measure demographic accentuation (Section 4.3) and report persona variants that achieve strong behavioral alignment while reducing demographic clustering.
@@
No sensitive protected health information or financial identifiers were collected. The anonymized dataset used for modeling contains only aggregated demographic categories and de-identified narrative text. Access to the synthetic dataset is restricted to the research team, and all analyses reported in this paper were conducted exclusively on the de-identified corpus. These procedures were designed to provide strong privacy protections while enabling rigorous sociotechnical research on persona-grounded AI systems.
Importantly, this work does not claim that AI models can replace or faithfully replicate individual human personas. Our evaluation framework is intentionally centered on patterns of behavioral similarity and aggregate tendencies, emphasizing structural alignment and averaged response behavior rather than individual-level substitution. The primary takeaway of this work is methodological: it demonstrates how persona construction choices, particularly the inclusion of sociopsychological structure beyond demographics—substantially affect behavioral realism, bias amplification, and downstream performance. We therefore position SCOPE as guidance on how to build better and more responsible personas, rather than as a claim of human replacement.
In regards to AI usage in the paper, we used AI-assisted tools solely for grammatical and stylistic corrections of our writing. All scientific content, analysis, and conclusions are the authors’ own.
@@ Appendix A Appendix
Refs: Park et al. (2023), Pew Research Center (2024), Davis and Smith (1999), Inglehart et al. (2014)
@@ A.7 External Data Formats: Nemotron Personas and SimBench Instances
Case 3: Demographics + Traits Demographics plus structured traits items and goal statements (Likert-style), emphasizing motivational priorities.
@@
Traits/Goals (snippet): “Helping people around me”=5; “Being rich”=2; “Equal opportunities”=5; “Taking risks/adventures”=3; “Secure surroundings”=4; “Thinking up new ideas/being creative”=4.
@@
Demographics + Traits/Goals
Case 4: Full Conditioning Combined structured persona: demographics + traits/goals + professional identity + personal identity narratives.
@@
Traits/Goals (snippet): “Environmental protection”=5; “Tradition”=2; “Achievement/recognition”=3; “Independence”=5.
@@
Demographics + Traits/Goals + Professional Identity + Personal Identity
@@
Case 7: Non-Demographic Persona Persona excludes demographics entirely; conditioning uses non-demographic signals (e.g., identity-only, traits-only, or traits+identity).
Example (traits+identity only):
@@
Traits/Goals (snippet): “Equality”=5; “Loyalty to friends”=5; “Tradition”=2; “Independence”=4; “Creativity”=3.
@@
Traits and/or Identity (no demographics)
@@
Case_2 Demographics + Narratives 0.642 37.74 0.011 8.298 0.656 38.56 0.012 8.791 0.609 33.58 -0.101 -76.79 0.616 27.58 -0.016 -12.245
Case_3 Demographics + Traits 0.656 38.60 0.100 75.499 0.650 39.60 0.055 41.287 0.600 36.71 0.012 9.087 0.614 22.56 -0.042 -31.64
@@
Case_6b AI (Given Demography + Narratives) 0.619 25.55 -0.034 -25.81 0.435 30.18 0.029 21.698 0.649 20.01 -0.041 -30.71 0.517 25.43 -0.065 -49.244
Case_6c AI (Given Demography + Traits) 0.606 27.86 0.016 12.249 0.462 30.15 0.009 6.797 0.681 21.32 -0.025 -19.097 0.496 27.30 -0.066 -50.171
@@
Case_6e AI (Given Traits alone) 0.588 41.65 -0.080 -71.43 0.592 41.45 -0.071 -63.02 0.544 40.49 -0.101 -90.04 0.577 40.37 -0.099 -87.90
Case_6f AI (Given Identity + Traits) 0.608 43.42 -0.046 -40.40 0.614 43.15 -0.038 -33.91 0.561 40.76 -0.079 -70.22 0.586 42.43 -0.082 -72.79
@@
Case_7b Traits Only 0.624 37.95 -0.094 -71.092 0.639 38.92 -0.088 -66.353 0.595 36.28 -0.121 -91.436 0.595 34.27 -0.115 -87.329
Case_7c Traits+Identity 0.658 39.68 -0.074 -56.354 0.662 40.29 -0.053 -40.485 0.612 37.86 -0.099 -74.563 0.636 33.89 -0.101 -76.371
@@
Case_2 Demographics + Narrative 0.609 35.40 -0.022 -0.207 0.615 31.80 0.035 0.322 0.572 20.90 -0.020 -0.190
Case_3 Demographics + Traits 0.602 35.10 -0.011 -0.107 0.591 29.10 0.050 0.469 0.553 19.10 -0.033 -0.304
@@
Case_6c AI (Given Demography + Traits) 0.560 25.76 -0.018 -13.811 0.564 23.22 -0.015 -11.300 0.521 15.79 -0.017 -13.183
Case_6d AI (Given Narrative alone) 0.528 36.73 -0.082 -72.564 0.532 33.10 -0.067 -59.371 0.492 22.51 -0.078 -69.266
Case_6e AI (Given Traits alone) 0.543 38.52 -0.097 -85.907 0.547 34.71 -0.079 -70.288 0.506 23.61 -0.092 -82.002
Case_6f AI (Given Identity + Traits) 0.561 40.15 -0.067 -59.763 0.566 36.19 -0.055 -48.897 0.523 24.61 -0.064 -57.046
Case_7a Narrative Only 0.584 34.62 -0.105 -79.570 0.589 31.20 -0.086 -65.103 0.544 21.22 -0.100 -75.953
Case_7b Traits Only 0.576 35.10 -0.115 -86.958 0.581 31.63 -0.094 -71.147 0.537 21.51 -0.110 -83.005
Case_7c Traits+Identity 0.608 36.70 -0.090 -68.138 0.612 33.07 -0.074 -55.749 0.566 22.49 -0.086 -65.040